OpenStudy (charan3):

The inner diameter of a cylindrical wooden pipe is 24 cm and its outer diameter is 28 cm. The pipe is 35cm long. Find the the mass of the pipe if 1 cmcube of wood has a mass of 0.6 gm.

OpenStudy (vijeya3):

well,all you need to do is find the volume of the material that would be used to make the cylinder.Next multiply your answer with the value 0.6 gram and you'll get your answer. here is the method- volume=pi*height(square of outer radius-square of inner radius) V=22/7*35(14*14-12*12) V=5270cc 1 cc mass=o.6 g so for 5720 cc=0.6*5720 3432 gm or 3.432 kg. I hope this is helpful:)

OpenStudy (charan3):

Thank u soo much even my sir also not able find answer